OpenStudy (anonymous):

simplifying the rational expression: 2x/(x-5)*(x-4)/(x+1) Can someone help me simplify this? I already know the answers but I'm having trouble explaining the steps to get there.

OpenStudy (anonymous):

@wiseman905

OpenStudy (anonymous):

Wow, I am only freshman. I don't think I can do that.

OpenStudy (anonymous):

Oh darn.

OpenStudy (anonymous):

hold on i need to find my old notes lol

OpenStudy (anonymous):

take your time. (not really, i need to get this done. xD)

OpenStudy (anonymous):

are you in intermediate algebra

OpenStudy (anonymous):

algebra 2. e.e i was unfortunate enough to get the new common core course. they literally give us 2 examples, don't explain it, and then tell us to try and figure it out. -____-

OpenStudy (anonymous):

im in algebra 2, im a sophomore let mehelp

OpenStudy (anonymous):

im a junior and i still dont remember this

OpenStudy (anonymous):

here look at this, if it doesnt help let me know http://www.purplemath.com/modules/rtnldefs2.htm

OpenStudy (anonymous):

I already have the answer, and somewhat the steps, can you help me explain it in numbers?

OpenStudy (anonymous):

2x/(x-5)*(x-4)/(x+1) = 2x(x-4)/(x+1)(x-5) First, we factor all the numerators and denominators. Then, we simplify. We do this by duplicating all the factors. Finally, we multiply the numerators and denominators together. I need to put the numbers after each step..

OpenStudy (anonymous):

ok gimme one sec

OpenStudy (anonymous):

ok gray im going to leave you to solve this one ok?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

oh my.. its been forever >~<

OpenStudy (anonymous):

\[\frac{ 2x }{ x-5 } \times \frac{ x-4 }{ x+1 }\]

OpenStudy (anonymous):

How do I factor the numerators and denominators?

OpenStudy (anonymous):

Do you need to oversimplify it?

OpenStudy (anonymous):

No, my solution is: 2x(x-4)/(x+1)(x-5) my teacher said my solution was perfect.. I just had to explain it xD

OpenStudy (anonymous):

Why can't anyone solve this? D:

ganeshie8 (ganeshie8):

\(\frac{ 2x }{ x-5 } \times \frac{ x-4 }{ x+1 } \) Step 1 : Factor The given rational expressions are in factored form already. So nothing to do here :- \(\large \frac{ 2x }{ x-5 } \times \frac{ x-4 }{ x+1 }\) Step 2 : Simplify Since there are no common factors that my be cancelled in both numerators and denominators, the rational expression remains same :- \(\large \frac{ 2x }{ x-5 } \times \frac{ x-4 }{ x+1 }\) Step 3 : Multiply Multiply the fractions by multiplying the remaining numerators and denominators.Keep the expression in factored form :- \(\large \frac{ 2x(x-4) }{ (x-5)(x+1) } \)

OpenStudy (anonymous):

Thank you so much!

ganeshie8 (ganeshie8):

your teacher wants u justify it step by step

ganeshie8 (ganeshie8):

np... u wlc :)